# Quantum Algorithms in a Superposition of Spacetimes

### v2

@@ 7 $\mathsf{GapCVP}_{O\left(n^{2}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
Soundness of the reduction.
We show that the reduction is correct, that is, that when $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ then the output distributions of the circuits, denoted $\mathcal{D}\left(C^{0}\right)$, $\mathcal{D}\left(C^{1}\right)$ respectively, are statistically close, and when the input is in $\prod_{\text{NO}}$, then the output distributions are statistically far.
We start with soundness: In case $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{NO}}$, then $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)>d\cdot g(n)$. Recall $\mathbf{e}\in\mathbb{Z}^{n}_{\left(-\kappa\cdot 2^{\beta},\kappa\cdot 2^{\beta}\right)}$ and thus $\left\lVert\mathbf{e}\right\rVert\leq\sqrt{n}\cdot\kappa\cdot 2^{\beta}$. Next, due to $\beta:=\lfloor\log_{2}\left(B\right)\rfloor$ we have
$2^{\beta}\leq B:=\frac{g(n)\cdot d}{2\cdot\kappa\cdot\sqrt{n}}\enspace,$
which implies $\left\lVert\mathbf{e}\right\rVert\leq\frac{g(n)\cdot d}{2}$. It follows that for any $\mathbf{v}\in\mathcal{L}_{\mathbf{B}}$, $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{v}+\mathbf{e}\right)\leq\frac{g(n)\cdot d}{2}$. Also, since $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)>d\cdot g(n)$ and $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{v}\right)=0$ then
$\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{v}+\mathbf{e}+\mathbf{t}\right)=\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{e}+\mathbf{t}\right)\geq\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)-\left\lVert\mathbf{e}\right\rVert>d\cdot g(n)-\frac{g(n)\cdot d}{2}=\frac{g(n)\cdot d}{2}\enspace.$
It follows that the output distributions $\mathcal{D}\left(C^{0}\right)$, $\mathcal{D}\left(C^{1}\right)$ of the circuits don’t intersect at all, that is, $\left\lVert\mathcal{D}\left(C^{0}\right)-\mathcal{D}\left(C^{1}\right)\right\rVert_{TV}=1$, as needed for soundness.
Completeness of the reduction, first part (reducing the analysis to analyzing total variation distance of the noise vector $\mathbf{e}$).
We now turn to the part where $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$, and we would like to give an upper bound on the total variation distance between the output distributions of the two circuits. We have,
$\left\lVert\mathcal{D}\left(C^{0}\right)-\mathcal{D}\left(C^{1}\right)\right\rVert_{TV}=\left\lVert\mathcal{D}\left(\mathbf{v}+\mathbf{e}\right)-\mathcal{D}\left(\mathbf{v}+\mathbf{e}+\mathbf{t}\right)\right\rVert_{TV}\enspace.$
We would like to use the fact that adding a lattice vector to a uniform distribution over that lattice, does not change the distribution. However, the lattice is infinite and we cannot sample a uniformly random vector over the entire lattice. Formally, since $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ then $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$, thus there exist $\mathbf{s}\in\mathcal{L}_{\mathbf{B}}$ and $\mathbf{e}_{0}\in\mathbb{Z}^{n}$ such that $\mathbf{t}=\mathbf{s}+\mathbf{e}_{0}$ and $\left\lVert\mathbf{e}_{0}\right\rVert\leq d$. It follows that,
$\left\lVert\mathcal{D}\left(\mathbf{v}+\mathbf{e}\right)-\mathcal{D}\left(\mathbf{v}+\mathbf{e}+\mathbf{t}\right)\right\rVert_{TV}$
(triangle $\underset{(\text{triangle inequality})}{\leq}\left\lVert\mathcal{D}\left(\mathbf{v}+\mathbf{e}\right)-\mathcal{D}\left(\mathbf{v}-\mathbf{s}+\mathbf{e}+\mathbf{t}\right)\right\rVert_{TV}+\left\lVert\mathcal{D}\left(\mathbf{v}+\mathbf{e}+\mathbf{t}\right)-\mathcal{D}\left(\mathbf{v}-\mathbf{s}+\mathbf{e}+\mathbf{t}\right)\right\rVert_{TV}$
$=\left\lVert\mathcal{D}\left(\left(\mathbf{v}\right)+\mathbf{e}\right)-\mathcal{D}\left(\left(\mathbf{v}\right)+\mathbf{e}+\mathbf{e}_{0}\right)\right\rVert_{TV}+\left\lVert\mathcal{D}\left(\mathbf{v}+\left(\mathbf{e}+\mathbf{t}\right)\right)-\mathcal{D}\left(\mathbf{v}-\mathbf{s}+\left(\mathbf{e}+\mathbf{t}\right)\right)\right\rVert_{TV}$
(adding variables can only reduce TV $\underset{(\text{adding variables can only reduce TV distance})}{\leq}\left\lVert\mathcal{D}\left(\mathbf{e}\right)-\mathcal{D}\left(\mathbf{e}+\mathbf{e}_{0}\right)\right\rVert_{TV}+\left\lVert\mathcal{D}\left(\mathbf{v}\right)-\mathcal{D}\left(\mathbf{v}-\mathbf{s}\right)\right\rVert_{TV}\enspace.$
We use the fact that we took the bound $M$ (which is effectively the sample space of our coordinates for the vector $\mathbf{v}$) to be exponentially (in the input dimension $n$) larger than the vectors in $\mathbf{B}$ and $\mathbf{t}$. Formally, note that since the norm of $\mathbf{s}$ is $\leq\left\lVert\mathbf{t}\right\rVert+\left\lVert\mathbf{e}_{0}\right\rVert\leq\left\lVert\mathbf{t}\right\rVert+d$, then $\left\lVert\mathbf{s}\right\rVert\cdot 2^{n}\leq M$. Since the variable $\mathbf{v}$ has a uniformly random coordinates vector in ($\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$, the length $\mathbf{s}$ is tiny compared to $\mathbf{v}$ with overwhelming probability. It follows that the distribution $\mathbf{v}-\mathbf{s}$ has total variation distance bounded by $2^{-n}$ to $\mathbf{v}$, and the above sum of total variation distances is bounded by
$\left\lVert\mathcal{D}\left(\mathbf{e}\right)-\mathcal{D}\left(\mathbf{e}+\mathbf{e}_{0}\right)\right\rVert_{TV}+2^{-n}\enspace.$
It remains to bound the total variation distance between $\mathcal{D}\left(\mathbf{e}\right)$ and $\mathcal{D}\left(\mathbf{e}+\mathbf{e}_{0}\right)$ for any $\mathbf{e}_{0}$ such that $\left\lVert\mathbf{e}_{0}\right\rVert\leq d$.
Completeness of the reduction, second part (moving to discrete Gaussians).
We have,
$\left\lVert\mathcal{D}\left(\mathbf{e}\right)-\mathcal{D}\left(\mathbf{e}+\mathbf{e}_{0}\right)\right\rVert_{TV}\underset{(\text{triangle inequality})}{\leq}\left\lVert\mathcal{D}\left(\mathbf{e}\right)-\mathcal{D}\left(D^{n}_{\tilde{B}}\right)\right\rVert_{TV}$
$+\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(D^{n}_{\tilde{B}}+\mathbf{e}_{0}\right)\right\rVert_{TV}$
$+\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}+\mathbf{e}_{0}\right)-\mathcal{D}\left(\mathbf{e}+\mathbf{e}_{0}\right)\right\rVert_{TV}\enspace,$
where $\tilde{B}:=\lceil\sqrt{\kappa}\cdot\tilde{\sigma}\rceil$, $\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}$, and the distribution $D^{n}_{\tilde{B}}$ is the $n$-dimensional discrete truncated Gaussian distribution, as defined in Definition 3.12. We use once again the fact that adding a numerical random variable to a pair of sums does not make it easier to distinguish the sums, and thus $\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}+\mathbf{e}_{0}\right)-\mathcal{D}\left(\mathbf{e}+\mathbf{e}_{0}\right)\right\rVert_{TV}\leq\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{e}\right)\right\rVert_{TV}$, and overall we have,
$\left\lVert\mathcal{D}\left(\mathbf{e}\right)-\mathcal{D}\left(\mathbf{e}+\mathbf{e}_{0}\right)\right\rVert_{TV}$
$\leq 2\cdot\left\lVert\mathcal{D}\left(\mathbf{e}\right)-\mathcal{D}\left(D^{n}_{\tilde{B}}\right)\right\rVert_{TV}+\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(D^{n}_{\tilde{B}}+\mathbf{e}_{0}\right)\right\rVert_{TV}\enspace.$
@@
where $\mathbf{e}_{1}$ is the distribution of the first coordinate of the vector $\mathbf{e}$ (which distributes identically for all of its coordinates $i\in[n]$). Recall that $\mathbf{e}_{1}$ is a random variable that’s the sum of $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-2^{\beta},2^{\beta}-1\right)}$ and $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$. By Lemma 7.1, there exists a positive absolute constant $c_{0}\in\mathbb{N}$ such that $\left\lVert\mathcal{D}\left(\mathbf{e}_{1}\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}\leq c_{0}\cdot\frac{1}{\sqrt{\kappa}}$.
@@
Completeness of the reduction, third part (Discrete Gaussians with close centers are close in total variation distance).
@@
$2^{-n}+\frac{c_{0}\cdot 2\cdot n}{\sqrt{\kappa}}+\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(D^{n}_{\tilde{B}}+\mathbf{e}_{0}\right)\right\rVert_{TV}\enspace,$
and it remains to give an upper bound for the last summand. Lemma 7.2 tells us that,
$\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(D^{n}_{\tilde{B}}+\mathbf{e}_{0}\right)\right\rVert_{TV}\leq\sqrt{1-e^{-\frac{\pi\cdot\left(\left\lVert\mathbf{e}_{0}\right\rVert^{2}+2\cdot\left\lVert\mathbf{e}_{0}\right\rVert\cdot\sqrt{n}\cdot\tilde{B}\right)}{\tilde{B}^{2}}}}\enspace.$
Now, recall that $\left\lVert\mathbf{e}_{0}\right\rVert\leq d$ and $\tilde{B}:=\lceil\sqrt{\kappa}\cdot\tilde{\sigma}\rceil$ for $\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}$. By further remembering that $2^{\beta}=\Theta\left(B\right)$, we get $\tilde{B}=\Theta\left(\frac{g(n)\cdot d}{\sqrt{n}\sqrt{\kappa}}\right)$. This means that there exists a positive absolute constant $c_{1}\in\mathbb{R}_{>0}$ such that $\tilde{B}\geq\frac{c_{1}\cdot g(n)\cdot d}{\sqrt{n}\sqrt{\kappa}}$. We get the following lower bound,

### v3

@@ 7 $\mathsf{GapCVP}_{O\left(n\sqrt{n}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
Sequentially invertible approximation of the discrete Gaussian distribution.
Recall the parameter $d_{\left(\mathbf{e},D_{\tilde{B}}^{n}\right)}$ from Equation 11: For both, the proofs of soundness and completeness of the reduction, we need to bound the absolute value of $d_{\left(\mathbf{e},D_{\tilde{B}}^{n}\right)}$. By the definition of $d_{\left(\mathbf{e},D_{\tilde{B}}^{n}\right)}$ it will be sufficient to bound $\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{e}\right)\right\rVert_{TV}$. Observe that the distance $\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{e}\right)\right\rVert_{TV}$ is essentially the quality of our (sequentially invertible) approximation of the truncated discrete Gaussian distribution.
@@
where $\mathbf{e}_{1}$ is the distribution of the first coordinate of the vector $\mathbf{e}$ (which distributes identically for all of its coordinates $i\in[n]$). By Lemma 7.1, there exists a positive absolute constant $c_{0}\in\mathbb{N}$ such that $\left\lVert\mathcal{D}\left(\mathbf{e}_{1}\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}\leq c_{0}\cdot\frac{1}{\sqrt{\kappa}}$. In conclusion, $|d_{\left(\mathbf{e},D_{\tilde{B}}^{n}\right)}|\leq\frac{c_{0}\cdot n}{\sqrt{\kappa}}\leq\frac{1}{8}$, thus $|2\cdot d_{\left(\mathbf{e},D_{\tilde{B}}^{n}\right)}|\leq\frac{1}{4}$.
Soundness of the reduction.
In case $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{NO}}$, then $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)>d\cdot g(n)$, and we should give a lower bound for
$\left\lVert\mathcal{D}\left(C^{0}\right)-\mathcal{D}\left(C^{1}\right)\right\rVert_{TV}$
$=\left\lVert\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}+\mathbf{t}\right)\right\rVert_{TV}+2\cdot d_{\left(\mathbf{e},D^{n}_{\tilde{B}}\right)}$
$\geq\left\lVert\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}+\mathbf{t}\right)\right\rVert_{TV}-\frac{1}{4}\enspace.$
Recall that,
$\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}<\sqrt{\frac{8\cdot 2^{2\beta}}{12}}<2^{\beta}\underset{(\beta:=\lfloor\log_{2}\left(B\right)\rfloor)}{\leq}B\enspace.$
Also,
$\tilde{B}:=\lceil\sqrt{\kappa}\cdot\tilde{\sigma}\rceil<\lceil\sqrt{\kappa}\cdot B\rceil:=\lceil\sqrt{\kappa}\cdot\frac{g(n)\cdot d}{c_{\mathcal{N}}\cdot\sqrt{n}\cdot\sqrt{\kappa}}\rceil=\lceil\frac{g(n)\cdot d}{c_{\mathcal{N}}\cdot\sqrt{n}}\rceil\underset{\left(\frac{g(n)\cdot d}{c_{\mathcal{N}}\cdot\sqrt{n}}>1\right)}{<}\frac{2\cdot g(n)\cdot d}{c_{\mathcal{N}}\cdot\sqrt{n}}\enspace.$
Now, Fact 3.4 tells us that with probability at least $1-e^{-c_{D}\cdot n}$ (for some absolute constant $c_{D}\in\mathbb{R}_{>0}$), the norm of the variable $\mathbf{u}\leftarrow D_{\tilde{B}}^{n}$ is bounded by $2\cdot\sqrt{n}\cdot\tilde{B}$, which in turn is bounded by $\frac{g(n)\cdot d}{2}$ due to $c_{\mathcal{N}}:=8$. It follows that for a sample $\mathbf{u}\leftarrow D_{\tilde{B}}^{n}$, with probability $\geq 1-e^{-c_{D}\cdot n}$ we have $\left\lVert\mathbf{u}\right\rVert\leq\frac{g(n)\cdot d}{2}$.
Now, consider the output distributions $\mathcal{D}\left(\mathbf{v}+\mathbf{u}\right)$, $\mathcal{D}\left(\mathbf{v}+\mathbf{u}+\mathbf{t}\right)$, for the case $\left\lVert\mathbf{u}\right\rVert\leq\frac{g(n)\cdot d}{2}$. For any $\mathbf{v}\in\mathcal{L}_{\mathbf{B}}$, $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{v}+\mathbf{u}\right)\leq\frac{g(n)\cdot d}{2}$. Also, since $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)>d\cdot g(n)$ and $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{v}\right)=0$ then
$\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{v}+\mathbf{u}+\mathbf{t}\right)=\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{u}+\mathbf{t}\right)\geq\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)-\left\lVert\mathbf{u}\right\rVert>d\cdot g(n)-\frac{g(n)\cdot d}{2}=\frac{g(n)\cdot d}{2}\enspace.$
It follows that if $\left\lVert\mathbf{u}\right\rVert\leq\frac{g(n)\cdot d}{2}$ then the distributions $\mathcal{D}\left(\mathbf{v}+\mathbf{u}\right)$, $\mathcal{D}\left(\mathbf{v}+\mathbf{u}+\mathbf{t}\right)$ don’t intersect, and since this happens with probability $\geq 1-e^{-c_{D}\cdot n}$, then
$\left\lVert\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}+\mathbf{t}\right)\right\rVert_{TV}\geq 1-e^{-c_{D}\cdot n}\enspace,$
which finally implies the reduction’s soundness, i.e., if $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{NO}}$ then
$\left\lVert\mathcal{D}\left(C^{0}\right)-\mathcal{D}\left(C^{1}\right)\right\rVert_{TV}\geq\frac{3}{4}-e^{-c_{D}\cdot n}\geq\frac{3}{4}-e^{-\Omega\left(n\right)}\enspace.$
Completeness of the reduction, first part (reducing the analysis to analyzing total variation distance of the noise vector $\mathbf{u}$).
In case $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$, then $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$, and we should give an upper bound for
$\left\lVert\mathcal{D}\left(C^{0}\right)-\mathcal{D}\left(C^{1}\right)\right\rVert_{TV}=\left\lVert\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}+\mathbf{t}\right)\right\rVert_{TV}+2\cdot d_{\left(\mathbf{e},D^{n}_{\tilde{B}}\right)}$
$\leq\left\lVert\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}+\mathbf{t}\right)\right\rVert_{TV}+\frac{1}{4}\enspace.$
Define $\mathbf{u}\leftarrow D^{n}_{\tilde{B}}$ as a sample from the distribution $D^{n}_{\tilde{B}}$. We would like to use the fact that adding a lattice vector to a uniform distribution over that lattice, does not change the distribution. However, the lattice is infinite and we cannot sample a uniformly random vector over the entire lattice (in finite time). Formally, since $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$, thus there exist $\mathbf{s}\in\mathcal{L}_{\mathbf{B}}$ and $\mathbf{e}_{0}\in\mathbb{Z}^{n}$ such that $\mathbf{t}=\mathbf{s}+\mathbf{e}_{0}$ and $\left\lVert\mathbf{e}_{0}\right\rVert\leq d$. It follows that,
$\left\lVert\mathcal{D}\left(\mathbf{v}+\mathbf{u}\right)-\mathcal{D}\left(\mathbf{v}+\mathbf{u}+\mathbf{t}\right)\right\rVert_{TV}$
(triangle $\underset{(\text{triangle inequality})}{\leq}\left\lVert\mathcal{D}\left(\mathbf{v}+\mathbf{u}\right)-\mathcal{D}\left(\mathbf{v}-\mathbf{s}+\mathbf{u}+\mathbf{t}\right)\right\rVert_{TV}+\left\lVert\mathcal{D}\left(\mathbf{v}+\mathbf{u}+\mathbf{t}\right)-\mathcal{D}\left(\mathbf{v}-\mathbf{s}+\mathbf{u}+\mathbf{t}\right)\right\rVert_{TV}$
$=\left\lVert\mathcal{D}\left(\left(\mathbf{v}\right)+\mathbf{u}\right)-\mathcal{D}\left(\left(\mathbf{v}\right)+\mathbf{u}+\mathbf{e}_{0}\right)\right\rVert_{TV}+\left\lVert\mathcal{D}\left(\mathbf{v}+\left(\mathbf{u}+\mathbf{t}\right)\right)-\mathcal{D}\left(\mathbf{v}-\mathbf{s}+\left(\mathbf{u}+\mathbf{t}\right)\right)\right\rVert_{TV}$
(summing more independent variables can only reduce TV $\underset{(\text{summing more independent variables can only reduce TV distance})}{\leq}\left\lVert\mathcal{D}\left(\mathbf{u}\right)-\mathcal{D}\left(\mathbf{u}+\mathbf{e}_{0}\right)\right\rVert_{TV}+\left\lVert\mathcal{D}\left(\mathbf{v}\right)-\mathcal{D}\left(\mathbf{v}-\mathbf{s}\right)\right\rVert_{TV}\enspace.$
We now use the fact that we took the bound $M$ (which is effectively the sample space of our coordinates for the vector $\mathbf{v}$) to be exponentially (in the input dimension $n$) larger than the vectors in $\mathbf{B}$ and $\mathbf{t}$. Formally, note that since the norm of $\mathbf{s}$ is $\leq\left\lVert\mathbf{t}\right\rVert+\left\lVert\mathbf{e}_{0}\right\rVert\leq\left\lVert\mathbf{t}\right\rVert+d$, then $\left\lVert\mathbf{s}\right\rVert\cdot 2^{n}\leq M$. Since the variable $\mathbf{v}$ has a uniformly random coordinates vector in ($\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$, the length $\mathbf{s}$ is tiny compared to $\mathbf{v}$ with overwhelming probability. It follows that the distribution $\mathbf{v}-\mathbf{s}$ has total variation distance bounded by $2^{-n}$ to $\mathbf{v}$, and the above sum of total variation distances is bounded by
$\left\lVert\mathcal{D}\left(\mathbf{u}\right)-\mathcal{D}\left(\mathbf{u}+\mathbf{e}_{0}\right)\right\rVert_{TV}+2^{-n}\enspace.$
It remains to bound the total variation distance between $\mathcal{D}\left(\mathbf{u}\right)$ and $\mathcal{D}\left(\mathbf{u}+\mathbf{e}_{0}\right)$ for any $\mathbf{e}_{0}$ such that $\left\lVert\mathbf{e}_{0}\right\rVert\leq d$.
@@
Completeness of the reduction, second part (Discrete Gaussians with close centers are close in total variation distance).
@@
$\left\lVert\mathcal{D}\left(\mathbf{u}\right)-\mathcal{D}\left(\mathbf{u}+\mathbf{e}_{0}\right)\right\rVert_{TV}+\frac{1}{4}+2^{-n}\enspace.$
Recall that $\mathbf{u}\leftarrow D^{n}_{\tilde{B}}$, and Lemma 7.2 tells us that
$\left\lVert\mathcal{D}\left(\mathbf{u}\right)-\mathcal{D}\left(\mathbf{u}+\mathbf{e}_{0}\right)\right\rVert_{TV}\leq\sqrt{1-e^{-\frac{\pi\cdot\left(\left\lVert\mathbf{e}_{0}\right\rVert^{2}+2\cdot\left\lVert\mathbf{e}_{0}\right\rVert\cdot\sqrt{n}\cdot\tilde{B}\right)}{\tilde{B}^{2}}}}\enspace.$
We have $\left\lVert\mathbf{e}_{0}\right\rVert\leq d$, and we previously set,
  • $g(n):=c_{g}\cdot n\sqrt{n}$.
  • $B:=\frac{g(n)\cdot d}{c_{\mathcal{N}}\cdot\sqrt{n}\cdot\sqrt{\kappa}}$, where $c_{\mathcal{N}}:=8$.
  • $\beta:=\lfloor\log_{2}\left(B\right)\rfloor$.
  • $\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}$.
  • $\tilde{B}:=\lceil\sqrt{\kappa}\cdot\tilde{\sigma}\rceil$.
The above implies
$\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}>\sqrt{\frac{4\cdot 2^{2\beta}}{12}}=2^{\beta}\cdot\frac{1}{\sqrt{3}}>\frac{2^{\beta}}{2}:=\frac{2^{\lfloor\log_{2}\left(B\right)\rfloor}}{2}\geq\frac{B}{4}:=\frac{g(n)\cdot d}{32\cdot\sqrt{n}\cdot\sqrt{\kappa}}\enspace,$
which in turn tells us that $\tilde{B}>\frac{g(n)\cdot d}{32\cdot\sqrt{n}}$. We now use this lower bound for $\tilde{B}$ for giving a lower bound for the exponent above: